OpenStudy (anonymous):

What is the sum of the geometric series 2^0+2^1+2^2+2^3+2^3+2^4+--+2^9? A: 1,023 B: 511 C: 2,407 D: 1,012

hartnn (hartnn):

do you know the formula of sum of geometric sequence

hartnn (hartnn):

ever seen this formula ? \(\large S_n= a_1 \dfrac{r^n-1}{r-1}\) ?

OpenStudy (anonymous):

Yes

hartnn (hartnn):

can you find a1 ? r ? n ?

hartnn (hartnn):

or know what they are ?

hartnn (hartnn):

a1 = 1st term = .... ? r = common ratio = 2^1/2^0 = 2^2/2^1 =... ? can you find these ?

OpenStudy (anonymous):

2^0 is the. 1st term

hartnn (hartnn):

correct! :) a1 = 2^0 =1 what about r ?

OpenStudy (anonymous):

Like 2^0, 2^1,2^3

hartnn (hartnn):

r = common ratio it is the number that you multiply to current term to get next term so, here r = 2^1/2^0 = ... ?

OpenStudy (anonymous):

2?

hartnn (hartnn):

correct! r= 2 :) from 2^0 to 2^9 , there are 10 terms, right ? so, n=10 any doubts till here ?

OpenStudy (anonymous):

Nope

hartnn (hartnn):

good,so plug these in that formula and simplify ... can you or need help with that ?

OpenStudy (anonymous):

I got it and the answer was A, thank you

hartnn (hartnn):

thats correct! S = r^10-1 = 1024 -1 = 1023 :) good :)

hartnn (hartnn):

and since you're new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

Thank you

hartnn (hartnn):

welcome ^_^